OpenStudy (anonymous):

Five added to twice a number is equal to 7 less than 4 times the number. Write the equation. 5+1^2=4x-7

OpenStudy (anonymous):

@johnweldon1993 @Hero @Marco,Phillip @undeadknight26 @ganeshie8 @e.mccormick @beccaboo333

OpenStudy (anonymous):

looks correct to me

OpenStudy (anonymous):

five added to twice a number. 5+ x^2 only thing i would fix there.

OpenStudy (anonymous):

Five added to twice a number is equal to 7 less than 4 times the number. What is the number? 3.25?

OpenStudy (anonymous):

if you work it out then yes you are correct

OpenStudy (anonymous):

The answer was actually 6. :)

OpenStudy (e.mccormick):

Twice a number = 2x

OpenStudy (e.mccormick):

5+2x=4x-7 5+7=4x-2x 12=2x 12/2=x 6=x